OpenStudy (jaydamarie1):

Which expression has a value of 74? (75−5)×1+4(75−5)×1+4 75−(5×1+4)75−(5×1+4) 75−5×(1+4)75−5×(1+4) (75−5)×(1+4)

OpenStudy (3mar):

May I help?

OpenStudy (jaydamarie1):

yes

OpenStudy (ivitt):

Do you still need some help?

OpenStudy (ivitt):

Or is 3mar going to assist you.

OpenStudy (jaydamarie1):

yes

OpenStudy (3mar):

Well, I am here.

OpenStudy (3mar):

What did you pick?

OpenStudy (jaydamarie1):

i pick c but i don't know if it is the right answer

OpenStudy (3mar):

so you need a check?

OpenStudy (jaydamarie1):

yes

OpenStudy (3mar):

so let's test it

OpenStudy (jaydamarie1):

ok

OpenStudy (3mar):

faster

OpenStudy (3mar):

1+4?

OpenStudy (jaydamarie1):

5

OpenStudy (3mar):

*5

OpenStudy (jaydamarie1):

25

OpenStudy (3mar):

*75

OpenStudy (jaydamarie1):

1875

OpenStudy (3mar):

good 75-1875?

OpenStudy (jaydamarie1):

1800

OpenStudy (3mar):

-5(1+4)?

OpenStudy (3mar):

-1800-5(1+4)=-1800-5(5)=-1800-25=??

OpenStudy (3mar):

@jaydamarie1

OpenStudy (jaydamarie1):

sry

OpenStudy (3mar):

Thank you for the medal!

OpenStudy (3mar):

Thank you for learning!!